OpenStudy (anonymous):

Every evening jenna emptys her pockets and puts her change in a jar. At the end of the week she counts her money. One week she had 40 coins, all of them dimes and quarters. When she added them up, she had a total of $7.75. How many quarters did jenna have?

OpenStudy (anonymous):

d=number of dimes q=number of quarters

OpenStudy (kirbykirby):

1st equation: d + q = 40 , does this make sense?

OpenStudy (anonymous):

Yea

OpenStudy (kirbykirby):

2nd equation: one dime = $0.10 one quarter = $0.25 so, the dollar $ amount in dimes you have is 0.10x and the dollar amount in quarters you have is 0.25y So, the total amount in quarter and dimes is 0.10x + 0.25y = 7.75

OpenStudy (anonymous):

Ok i see

OpenStudy (anonymous):

How do we find out how many quarters we had??

OpenStudy (kirbykirby):

(sorry I accidentally replaced d and q with x and y... so you should have 0.10d + 0.25q = 7.75) The number of quarters is q, But you have 2 equations and 2 unknowns. So you can do like in the previous problem... and solve for d and q: d + q = 40 0.10q + 0.25q = 7.75

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

solve for q in both equations: equation 1: \[d + q = 40\\ d + q - d = 40 - d\\ q = 40 - d\] second equation: \[0.10d + 0.25q = 7.75\\ 0.10d + 0.25q - 0.10d = 7.75 - 0.10d\\ 0.25q = 7.75 - 0.10d\] \[\frac{0.25q}{0.25} = \frac{7.75 - 0.10d}{0.25}\\ \, \\ q = \frac{7.75-0.10d}{0.25}\] Now equate both equations: \[40 - d = \frac{7.75-0.10d}{0.25} \\ \, \\ 0.25(40-d)=7.75-0.10d\ \, \\ 10-0.25d = 7.75 - 0.10d\\ -0.25d+0.10d=7.75-10\\ -0.15d=-2.25\\ d=\frac{-2.25}{-0.15}\\ d=15 \] Now back to the first equation, since you know d: \[q + d = 40\\ q + 15 = 40\\ q = 40 - 15\\ q = 25\]